# Cascade Speculative Drafting for Even Faster LLM Inference

### 21 Dec 2023 v2

@@ 3.3 Max-gram for Better Statistical Drafting
A statistical language model is special to speculative decoding, as it incurs almost no cost when drafting for its target model. In our pursuit of a more effective statistical language model, we noticed a general pattern: in language model generation, some words and phrases from the input query frequently reappear in the generated content. Based on this observation, we designed the Max-Gram (MaG) algorithm. It greedily identifies maximal matches between the initial input (or existing generation) and tokens from the end of the generation. In cases where there is no match, we resort to a bigram model based on the probability distribution of Wikipedia (chosen to maintain the generality of our experiment). Surprisingly, despite its simplicity, this straightforward and general algorithm has proven highly effective in aligning with neural language model outputs. We include a GPU-friendly version of the Max-Gram algorithm in Appendix A.
@@ 3.4 Algorithm
Combining the horizontal and vertical cascades, the algorithm is presented in Algorithm 1. At its center, there is a for-loop that performs the horizontal cascade, while the recursive call forms the vertical cascade. Notably, the last element of the draft models is the MaG model.
The algorithm requires a two-dimensional hyperparameter, $K_{nn}$, with each row serving as the stop criteria for a layer of recursive calls. For simplicity, we assume the leniency parameter $l$ is universal for the algorithm, except when the target model is under review; thus, the algorithm can benefit from the speedup of leniency without altering the output distribution.
4 Theoretical Analysis
In this section, we present some theoretical analyses of the proposed Cascade Speculative Drafting algorithm.
4.1 Expected Walltime Improvement
To assist our theoretical analysts, we establish the notions. Let $\mathcal{M}_{t}$ be the target model, $\{\mathcal{M}_{i}\}$ be the draft models assisting generation with $\mathcal{M}_{i}$ being the draft model generating the $i$-th token. In the simpler case of the speculative decoding, $\mathcal{M}_{i}=\mathcal{M}_{d}$ for any $i$. Let $x$ be the input to the model at a single run, $\mathcal{M}_{t}(x)$ and $\mathcal{M}_{i}(x)$ are then the output probability distribution with input $x$.
Expected acceptance rate $\alpha=\mathbf{E}(1-\frac{1}{2}|\mathcal{M}_{t}(x)-\mathcal{M}_{i}(x)|)$ is the probability of draft generation by $\{\mathcal{M}_{i}\}$ being accepted by target model $\mathcal{M}_{t}$ (Leviathan et al., 2023).
Cost coefficient $c_{i}$ is the ratio of time for a single run of draft model $\mathcal{M}_{i}$ over target model $\mathcal{M}_{t}$.
Expected walltime improvement factor (EWIF) is the expected time improvement of an algorithm based on. We provide the formula for EWIF in an extended setting when multiple draft models are involved.
@@ 4.1 Expected Walltime Improvement
The expected walltime improvement factor (EWIF) by Algorithm 1 is $T(k,{\alpha_{1},...,\alpha_{k}},{c_{1},...,c_{k}})=\frac{\sum_{i=0}^{k}\prod_{j=1}^{i}\alpha_{j}}{1+\sum_{i=1}^{k}c_{i}}$.
Proof.
Since every token will only be considered by the target model if all tokens before are accepted, a simple recursion leads to the probability of acceptance of the $i$-th token being $\prod_{j=1}^{i-1}\alpha_{j}$, and the $\prod_{j=1}^{0}\alpha_{j}=1$ representing the token generated by the target model. In the right-hand side, the numerator is the expected number of tokens generated, and the denominator is the time costs for each step of such a system, so the right-hand side is the expected efficiency improvement. ∎
We can immediately achieve a formula for speculative decoding by requiring all the draft models to be the same.
@@
When $q_{1}=q_{2}=...=q_{k}$, let $\alpha_{i}=\alpha$ and $c_{i}=c$ for any $i$, the EWIF is $\frac{1-\alpha^{k+1}}{(1-\alpha)(c\cdot k+1)}$.
@@
Using the information provided by Leviathan et al. (2023), we calculat a simulated EWIF under the assumption that the event of acceptance by the target model is a Bernoulli trial. The results, shown in Table 1, indicate that speculative sampling with a horizontal cascade achieved better EWIF than vanilla speculative sampling under this assumption.
4.2 Analysis of Model Interpolation
The setting of cascade drafting can be simply extended using model interpolation that involves choosing between two models using randomness. Given two draft systems $d_{1}$ and $d_{2}$ (They can be two models generating autoregressive or two speculative decoding systems as used in vertical cascade) and a probability function $f$, an interpolation $d_{inter}$ is defined as $f(i)d_{1}+(1-f(i))d_{2}$. In other words, with probability $f(i)$, $d_{inter}$ is $d_{1}$, and with probability $1-f(i)$, $d_{inter}$ is $d_{2}$. As a result, $\alpha_{i}=f(i)\alpha_{d1}+(1-f(i))\alpha_{d2}=\alpha_{d2}+f(i)(\alpha_{d1}-\alpha_{d2})$ and $c_{i}=f(i)c_{d1}+(1-f(i))c_{d2}=c_{2}+f(i)(c_{d1}-c_{d2})$. An interpolation system can provide a smooth transition between two systems; however, in theorem 4.4, we prove that the optimal can always be reached without interpolation.
Theorem 4.4.
Given a horizontal cascade system with the setting mentioned above. There exists a probability function $f$ that optimizes the system such that $f$ satisfies $f(i)\in\{0,1\}$ for any $i$. This means each of the tokens will be generated by either $d_{1}$ or $d_{2}$, and the optimal system is deterministic.
Proof of Theorem 4.4 is included in appendix B.
Dataset $\mathcal{M}_{d1}$ $\mathcal{M}_{d2}$ $k_{1}$ $k_{2}$ EWIF
@@ 5 Evaluation
In addition, a recent analysis suggests GPU speed can vary for the same GPU model with one being 1.5x faster than another GPU of the same model (Sinha et al., 2022), and even for the same GPU, the performance variation between different runs is unknown; therefore, theoretical walltime improvement up to 1.5x can be undermined by GPU variation, and methods without any improvement can be measured with an improvement up to 1.5x by running on a different GPU. The properties of such lack of standardization and stability make walltime unfit for measuring algorithmic improvement. To resolve this issue, we design a new metric to measure speculative decoding and its related algorithms which can reduce the impact of varying/unstable GPU performance.
Our proposed method, standardized walltime improvement (SWI), calculates the GPU times of the models, assuming that each run of a language model costs the same amount of time, an assumption made when inventing the speculative sampling algorithm (Chen et al., 2023). Given a speculative decoding system $S$ containing $k$ draft models $\mathcal{M}_{d1},...,\mathcal{M}_{dk}$, target model $\mathcal{M}_{t}$ solving a task $T$, and let $count(S,m,T)$ be the count of runs of model $m$ during the execution of task $T$ with system $S$, let $c_{i}$ be the expected cost coefficient being the time ratio between a single run of $\mathcal{M}_{i}$ and $\mathcal{M}_{t}$, and let $n$ be the total of tokens generated, the standardized walltime improvement is defined as
$\begin{split}SWI&(\mathcal{M}_{d1},...,\mathcal{M}_{dk},\mathcal{M}_{t},T)=\\ &\frac{n}{count(S,\mathcal{M}_{t},T)+\sum_{i=1}^{n}count(S,\mathcal{M}_{di},T)c_{i}},\end{split}$
which is simply the sum of GPU time by all models when $c_{i}$ is measured during a local experiment. The benefit of the $SWI$ evaluation is that $count(S,\mathcal{M}_{t},T)$ and $count(S,\mathcal{M}_{di},T)$ will be the same across different systems, so by using the same $c_{i}$ across different experiments the GPU performance variation will be eliminated; therefore, it provides a standardized, reproducible measurement.
During our experiment, we will use two sets of $\{c_{i}\}$, existing $\{c_{i}\}$ reported by Leviathan et al. (2023) and an heuristic $\{c_{i}\}$ being the ratio of number of tunable parameters between $\mathcal{M}_{di}$ and $\mathcal{M}_{t}$.
Dataset Algorithm $\{\mathcal{M}_{di}\}$ SWI (MS) $k_{11}$ $k_{21}$ $k_{12}$ $l$
@@
GSM8K S. Decoding base $3.38$ 10 - - -
GSM8K S. Decoding small $3.06$ 11 - - -
GSM8K C.S. Drafting base, mag $3.62$ 9 - - -
GSM8K C.S. Drafting small, mag $3.16$ 11 - - -
GSM8K C.S. Drafting base, small, mag 3.75 7 1 10 1.5
@@
MMLU S. Decoding base $3.97$ 13 - - -
MMLU S. Decoding small $4.12$ 19 - - -
MMLU C.S. Drafting base, mag $4.50$ 12 - - -
MMLU C.S. Drafting small, mag $4.36$ 17 - - -
MMLU C.S. Drafting base, small, mag 4.83 5 1 19 2
Dataset Algorithm $\{\mathcal{M}_{di}\}$ SWI (PW) $k_{11}$ $k_{21}$ $k_{12}$ $l$
@@
GSM8K S. Decoding base $2.99$ 8 - - -
GSM8K S. Decoding small $2.76$ 8 - - -
GSM8K C.S. Drafting base, mag $3.22$ 8 - - -
GSM8K C.S. Drafting small, mag $2.90$ 7 - - -
GSM8K C.S. Drafting base, small, mag 3.29 5 1 9 1.5
@@
MMLU S. Decoding base $3.42$ 10 - - -
MMLU S. Decoding small $3.51$ 11 - - -
MMLU C.S. Drafting base, mag $4.07$ 9 - - -
MMLU C.S. Drafting small, mag $3.97$ 12 - - -
MMLU C.S. Drafting base, small, mag 4.23 5 1 13 1.5
Table 2: The experiment results from datasets GSM8K and MMLU. BASE and SMALL refer to FLAN-T5-base and FLAN-T5-small. SWI (MS) is the standardized walltime improvement with the assumption that the latency of each run of a model is its number of parameters (model size). SWI (PW) is the SWI with the assumption that the latency of each run of a model is the time cost data reported from previous work (Leviathan et al., 2023). $k_{11}$, $k_{21}$, $k_{12}$, $l$ are the hyperparameters. $k_{1i}$ represents the step limitation target model and the draft models, $k_{21}$ is the step limitations between the first and second draft model, and $l$ is leniency as shown in algorithm 1. For speculative decoding, the $k_{11}$ is simply the $k$.
6 Experiment
@@ 6.1 Setup
We chose our target and draft models from the FLAN-T5 (Chung et al., 2022) family for our experiment, as there is a large variation in model sizes within the FLAN-T5 family (ranging from 77 million to 11 billion parameters). To align our experiment with current common usage, we do not perform fine-tuning, and the generation is conducted in a zero-shot manner. We use FLAN-T5-xxl as our target model, FLAN-T5-base and FLAN-T5-small as our reviewing draft models, and the Max-Gram algorithm as the generating draft model.
To ensure our experiments are fully reproducible, we perform sampling at temperature 00, which is essentially greedy decoding. And we do not observe any significant difference between sampling with temperature $1$ and greedy decoding in previous speculative decoding experiments (Leviathan et al., 2023).
Cascade Speculative Drafting Inference To reduce the number of hyperparameters to tune, we use MaG to generate 10 tokens at once, as it is rare for more than 10 tokens to be accepted. We do not use lenience when the reviewer is $\mathcal{M}_{t}$ to ensure the output distribution remains does not change. Additionally, we eliminate leniency between $\mathcal{M}_{d2}$ and MaG because the performance gap between them is large.
With these constraints, we are left with at most four hyperparameters: $k_{11}$, $k_{21}$, $k_{12}$, and $l$. For the CS Drafting step where the target model is the reviewer, $k_{11}$ and $k_{12}$ are used. $k_{21}$ and $l$ are used in the step where $\mathcal{M}_{d1}$ is the reviewer. However, there are still too many hyperparameters for us to exhaustively search for the optimal result; instead, we employ a halving grid search to find a set of hyperparameters.
Baseline We consider both speculative decoding and vanilla autoregressive generation as our baselines. For speculative decoding, we opt for FLAN-T5-small and FLAN-T5-base as the draft models, as they perform better than FLAN-T5-large, according to Leviathan et al. (2023) and our preliminary experiments. We test a wide range of $k$-values, from 2 to 30, to ensure that we find the best result for each model serving as our draft model to establish our baseline.

### 16 Feb 2024 v3

@@ 3.3 Max-Gram for Better Statistical Drafting
As both Vertical Cascade and Horizontal Cascade remark cascade toward faster draft models, a statistical language model, which is the basis of the cascade, becomes essential for the efficiency of both approaches. In our pursuit of a more effective statistical language model, we noticed a general pattern: in language model generation, some words and phrases from the input query frequently reappear in the generated content. In light of this observation, we designed the Max-Gram (MaG) algorithm. It greedily identifies maximal matches between the initial input (or existing generation) and tokens from the end of the generation. In cases where there is no match, we resort to a bigram model based on the probability distribution of Wikipedia (chosen to maintain the generality). Surprisingly, despite its simplicity, this straightforward and general algorithm has proven highly effective in aligning with neural language model outputs. We include a GPU-friendly version of the Max-Gram algorithm in Appendix A.
@@ 3.4 Algorithm
Combining the horizontal and vertical cascades, the algorithm of cascade speculative decoding is presented in Algorithm 1. At its center, the horizontal cascade is realized by the for loop, while the vertical cascade is implemented through recursive calls. Notably, the MaG model is incorporated as the smallest draft model to avoid autoregressive generation from a neural model. An example of CS Drafting is shown in Figure 1.
The algorithm requires an upper-triangular hyperparameter, $K_{nn}$, with each row serving as the stop criteria for a layer of recursive calls. For simplicity, we assume the lenience parameter $l$ is universal for the algorithm, except when the target model is under review; thus, the algorithm can benefit from the speedup of lenience without altering the output distribution.
4 Analysis
We begin with some notions. Let $\mathcal{M}_{t}$ be the target model, $\mathcal{M}_{d}$ be the draft model, and $k$ be the number of draft tokens generated per step.
Expected acceptance rate $\alpha(\mathcal{M}_{t},\mathcal{M}_{d})$ is the probability of draft generation by $\mathcal{M}_{d}$ being accepted by target model $\mathcal{M}_{t}$.
Cost coefficient $c(\mathcal{M}_{t},\mathcal{M}_{d})$ is the ratio of time for a single run of draft model $\mathcal{M}_{d}$ over target model $\mathcal{M}_{t}$.
Expected walltime improvement factor (EWIF) is the expected time improvement achieved by an algorithm under the i.i.d. assumption of token acceptance.
Despite the simple setting of EWIF, Leviathan et al. (2023) have demonstrated that it aligns with the experimental results in most instances. Therefore, our analysis will concentrate on EWIF.
4.1 Vertical Cascade
We analyze EWIF of vertical cascade using generating functions, a well-studied topic in combinatorial mathematics (West, 2021). The properties of generating functions are useful in the recursion and evaluation process making the our final expression simple.
We begin with the derivation of the probability generating function for speculative decoding.
@@ 4.1 Vertical Cascade
For speculative decoding between $\mathcal{M}_{t}$ and $\mathcal{M}_{d}$, let ${p_{i}}$ be the probability of generating $i$ tokens. The probability generating function of ${p_{i}}$ satisfies the following equation:
$\phi_{(\alpha,k)}(x)=1+(x-1)\frac{1-\alpha^{k+1}x^{k+1}}{(1-\alpha x)},$ (1)
where $\alpha=\alpha(\mathcal{M}_{t},\mathcal{M}_{d})$.
Proof in Appendix C.1.
@@
The EWIF of speculative decoding is $\frac{\phi_{(\alpha,k)}^{\prime}(1)}{(ck+1)}=\frac{1-\alpha^{k+1}}{(1-\alpha)(ck+1)}$.
We use the generating function to derive the EWIF of a vertical cascade and analyze the case involving two draft models, $\mathcal{M}_{d_{1}}$ and $\mathcal{M}_{d_{2}}$.
Theorem 4.3.
Assume $k$ to be the speculative decoding parameter between $\mathcal{M}_{d_{1}}$ and $\mathcal{M}_{d_{2}}$, and $n$ to be the number of steps $\mathcal{M}_{t}$ reviews. The EWIF by this system is
$\frac{1-\alpha\phi^{n}(\alpha)}{(1-\alpha)(1+nc_{d_{1}}+nkc_{d_{2}})},$ (2)
where $\phi(x)=\phi_{(\alpha(\mathcal{M}_{d_{1}},\mathcal{M}_{d_{2}}),k)}(x)$, $\alpha=\alpha(\mathcal{M}_{t},\mathcal{M}_{d})$, and $c_{d_{1}},c_{d_{2}}$ be $c(\mathcal{M}_{t},\mathcal{M}_{d_{1}}),c(\mathcal{M}_{t},\mathcal{M}_{d_{2}})$ respectively.
Corollary 4.4.
$\phi_{\alpha^{\prime},k}(\alpha)<\alpha$ for any $01>\alpha>0,1>\alpha^{\prime}>0,k>$ 0, so if $c_{d_{2}}<<1$, the EWIF of $\mathcal{M}_{d_{1}}$ and $\mathcal{M}_{d_{2}}$ is higher than EWIF of $\mathcal{M}_{d_{1}}$ alone.
Proof in Appendix C.2.
Therefore, with the statistical model having negligible cost (i.e., $c_{d_{2}}<<1$), it can almost always improve the efficiency of an SD system.
4.2 Horizontal Cascade
We also present an analysis of the walltime improvement offered by the horizontal cascade.
To assist the analysis, we establish the notions. Let $\mathcal{M}_{t}$ be the target model, $\{\mathcal{M}_{i}\}$ be the draft models assisting generation with $\mathcal{M}_{i}$ being the draft model generating the $i$-th token. In the simpler case of the speculative decoding, $\mathcal{M}_{i}=\mathcal{M}_{d}$ for any $i$. Let $x$ be the input to the model at a single run, $\mathcal{M}_{t}(x)$ and $\mathcal{M}_{i}(x)$ are then the output probability distribution with input $x$. To simplify notation, let $\alpha_{i}=\alpha(\mathcal{M}_{t}(x),\mathcal{M}_{i}(x))$ and $c_{i}=c(\mathcal{M}_{t}(x),\mathcal{M}_{i}(x))$.
Theorem 4.5.
The expected walltime improvement factor (EWIF) of the horizontal cascade is $T(k,{\alpha_{1},...,\alpha_{k}},{c_{1},...,c_{k}})=\frac{\sum_{i=0}^{k}\prod_{j=1}^{i}\alpha_{j}}{1+\sum_{i=1}^{k}c_{i}}$.
@@ 4.2 Horizontal Cascade
Using the information provided by Leviathan et al. (2023), we calculate a simulated EWIF under the assumption that the event of acceptance by the target model is a Bernoulli trial. The results, shown in Table 1, indicate that speculative sampling with a horizontal cascade achieved better EWIF than vanilla speculative sampling under this assumption.
Dataset $\mathcal{M}_{d_{1}}$ $\mathcal{M}_{d_{2}}$ $k_{1}$ $k_{2}$ EWIF
@@ 5.1 Evaluation Metric
In addition, a recent analysis suggests GPU speed can vary for the same GPU model with one being 1.5x faster than another GPU of the same model (Sinha et al., 2022); therefore, theoretical walltime improvement up to 1.5x can be undermined by GPU variation, and methods without any improvement can be measured with an improvement by running on a different GPU. The lack of standardization and stability makes walltime unfit for measuring algorithmic improvement. To resolve this issue, we design a new metric to measure speculative decoding and its related algorithms which can reduce the impact of varying/unstable GPU performance.
Our proposed method, standardized walltime improvement (SWI), calculates the GPU times of the models, assuming that each run of a language model costs the same amount of time, an assumption made when inventing the speculative sampling algorithm (Chen et al., 2023). Given a speculative decoding system $S$ containing $k$ draft models $\mathcal{M}_{d_{1}},...,\mathcal{M}_{d_{k}}$, target model $\mathcal{M}_{t}$ solving a task $T$, and let $count(S,m,T)$ be the count of runs of model $m$ during the execution of task $T$ with system $S$, let $c_{i}$ be the expected cost coefficient being the time ratio between a single run of $\mathcal{M}_{i}$ and $\mathcal{M}_{t}$, and let $n$ be the total of tokens generated, the standardized walltime improvement is defined as
$\begin{split}SWI&(\mathcal{M}_{d_{1}},...,\mathcal{M}_{d_{k}},\mathcal{M}_{t},T)=\\ &\frac{n}{count(S,\mathcal{M}_{t},T)+\sum_{i=1}^{n}count(S,\mathcal{M}_{d_{i}},T)c_{i}},\end{split}$
which is simply the sum of GPU time by all models when $c_{i}$ is measured during a local experiment. The benefit of the $SWI$ evaluation is that $count(S,\mathcal{M}_{t},T)$ and $count(S,\mathcal{M}_{d_{i}},T)$ will be the consistent across different systems, so by using the same $c_{i}$ across different experiments the GPU performance variation will be eliminated; therefore, it provides a standardized, reproducible measurement.
During our experiment, we will use two sets of $\{c_{i}\}$, existing $\{c_{i}\}$ reported by Leviathan et al. (2023) and an heuristic $\{c_{i}\}$ being the ratio of number of tunable parameters between $\mathcal{M}_{d_{i}}$ and $\mathcal{M}_{t}$.
Dataset Algorithm $\{\mathcal{M}_{d_{i}}\}$ Speedup (MS) $k_{11}$ $k_{12}$ $k_{22}$ $l$
@@
GSM8K S Decoding base $3.38$ 10 - - -
GSM8K S Decoding small $3.06$ 11 - - -
GSM8K CS Drafting base, mag $3.70$ 10 - - -
GSM8K CS Drafting small, mag $3.19$ 11 - - -
GSM8K CS Drafting base, small, mag 3.88 8 13 1 3
@@
MMLU S Decoding base $3.97$ 13 - - -
MMLU S Decoding small $4.12$ 19 - - -
MMLU CS Drafting base, mag $4.56$ 13 - - -
MMLU CS Drafting small, mag $4.39$ 14 - - -
MMLU CS Drafting base, small, mag 4.88 5 19 1 5
Dataset Algorithm $\{\mathcal{M}_{d_{i}}\}$ Speedup (PW) $k_{11}$ $k_{12}$ $k_{22}$ $l$
@@
GSM8K S Decoding base $2.99$ 8 - - -
GSM8K S Decoding small $2.76$ 8 - - -
GSM8K CS Drafting base, mag $3.27$ 9 - - -
GSM8K CS Drafting small, mag $2.82$ 11 - - -
GSM8K CS Drafting base, small, mag 3.43 5 9 1 3
@@
MMLU S Decoding base $3.42$ 10 - - -
MMLU S Decoding small $3.51$ 11 - - -
MMLU CS Drafting base, mag $4.21$ 6 - - -
MMLU CS Drafting small, mag $3.99$ 13 - - -
MMLU CS Drafting base, small, mag 4.32 5 8 1 5
Table 2: The experimental results on FLAN-T5. Speedup (MS) is the standardized walltime improvement with the assumption that the latency of each run of a model is its number of parameters (model size). Speedup (PW) is the SWI with the assumption that the latency of each run of a model is the time cost data reported from previous work (Leviathan et al., 2023). $k_{11}$, $k_{12}$, $k_{22}$, $l$ are the hyperparameters. $k_{11}$ and $k_{12}$ represent the step limitation target model and the draft models, $k_{22}$ is the step limitations between the first and second draft model, and $l$ is lenience as shown in algorithm 1. For speculative decoding, the $k_{11}$ is simply the $k$.
@@ 5.2 Setup
To ensure the generality of our findings, we perform experiments on both encoder-decoder and decoder-only models. For encoder-decoder models, we choose our target and draft models from the FLAN-T5 (Chung et al., 2022) family for our experiment, as there is a large variation in model sizes within the FLAN-T5 family (ranging from 77 million to 11 billion parameters). We use FLAN-T5-xxl as our target model, FLAN-T5-base and FLAN-T5-small as our reviewing draft models. For decoder-only models, we select LLAMA-2-chat-7B (Touvron et al., 2023) as the target model. Since there is no official release of a smaller model in LLAMA family, we use a pretrained 160M model with the same tokenizer as the reviewing draft model (Miao et al., 2024). In both cases, the Max-Gram algorithm is used as the generating draft model. To align our experiment with current common usage, we do not perform fine-tuning, and the generation is conducted in a zero-shot manner.
Since we do not observe any significant difference between sampling with temperature $1$ and greedy decoding in previous speculative decoding experiments (Leviathan et al., 2023), and to ensure our experiments are fully reproducible, we perform sampling at temperature 00, i.e., using greedy decoding by default.
Cascade Speculative Drafting Inference To reduce the number of hyperparameters to tune, we use MaG to generate 10 tokens at once, as it is rare for more than 10 tokens to be accepted. We do not use lenience when the reviewer is $\mathcal{M}_{t}$ to ensure the output distribution does not change. We also avoid lenience between MaG and its reviewer, since there is still a significant performance gap between MaG and a neural model.
With these constraints, we are left with at most four hyperparameters: $k_{11}$, $k_{12}$, $k_{22}$, and $l$. For the CS Drafting step where the target model is the reviewer, $k_{11}$ and $k_{12}$ are used. $k_{21}$ and $l$ are used in the step where $\mathcal{M}_{d_{1}}$ is the reviewer.
Baselines We consider both speculative decoding and vanilla autoregressive generation as our baselines. When performing speculative decoding with the target model as FLAN-T5-xxl, we opt for FLAN-T5-small and FLAN-T5-base as the draft models, as they perform better than FLAN-T5-large, according to Leviathan et al. (2023) and our preliminary experiments. When LLAMA-2-chat-7B is the target model, we use the 160M model as the draft model.
We test a wide range of $k$-values, from 2 to 30, to ensure that we find the best result for each model serving as our draft model to establish our baseline.